OpenStudy (anonymous):

Find all solutions to the equation in the interval [0,2pi). sin2x - sin4x = 0

OpenStudy (anonymous):

sin4x can be written : 2sin(2x)cos(2x) does it help ?

OpenStudy (anonymous):

kind of. i'm still a bit confused though. so what i have so far is 2sinxcosx - 2sin(2x)cos(2x) = 0

OpenStudy (anonymous):

would i find a gcf next and factor out the 2sinxcosx?

OpenStudy (anonymous):

i think it is better to write it : sin(2x) - 2sin(2x)cos(2x)=0 and now you can factor sin(2x) out

OpenStudy (anonymous):

sin(2x) * (1-2cos(2x)) = 0 you have to find the values: sin(2x) = 0 1-2cos(2x) = 0

OpenStudy (freckles):

I don't know if you are familiar with this identity or not but if you are you can use: \[\sin(x)-\sin(y)=2 \cos(\frac{x+y}{2})\sin(\frac{x-y}{2})\]

OpenStudy (freckles):

and if you don't @Coolsector has a cool style too :)

OpenStudy (anonymous):

@freckles i haven't learned about sum to product formulas yet so i'm not familiar with that. i just learned about multiple angle formulas and sum and difference formulas so that's what i'm trying to use but thanks though!

OpenStudy (anonymous):

and @Coolsector thank you that helped!

OpenStudy (anonymous):

yw